Stubbs Company uses the perpetual inventory method. On January 1, Year 1, Stubbs purchased 700 units of inventory that cost $5.00 each. On January 10, Year 1, the company purchased an additional 600 units of inventory that cost $3.75 each. If Stubbs uses a weighted average cost flow method and sells 1,000 units of inventory for $10.00 each, the amount of gross margin reported on the income statement will be: (Round your intermediate calculations to two decimal places.)

Multiple Choice

  • $5,746.

  • $5,580.

  • $8,125.

  • $7,250.
Answer #1

Total cost = (700*5)+(600*3.75) = 5750

Total units = 700+600 = 1300

Average cost per unit = 5750/1300 = 4.42 per unit

Gross margin = Sales - Cost of Goods Sold

= (1,000*10) - (1,000*4.42)

= 5580

Option B is the answer